OpenStudy (anonymous):

For each function, y varies directly with x. Find each constant variation. Then find the value of y when x= -0.3 1. y=2 when x= -1/2 2. y= 2/3 when x= 0.2 3. y=7 when x= 2 4. y=4 when x= -3

OpenStudy (anonymous):

can someone at least explain it to me or show me an example? :c

OpenStudy (aum):

"y varies directly with x" means y = constant * x or y = kx ----- (1) 1. when x = -1/2, y = 2 Put it in (1): 2 = k * (-1/2) multiply by 2: 4 = k * (-1) 4 = -k k = -4 y = -4x "find the value of y when x= -0.3" y = -4x = (-4) * (-0.3) = 1.2

OpenStudy (aum):

So for each problem, start with y = kx They give you one x and one y value. Put it in y = kx and find k. Then put x = -0.3 and find y.

OpenStudy (anonymous):

Why would i multiply it though; for the first one?;o

OpenStudy (anonymous):

The easiest way to do these type of problems is to set it up like a proportion

OpenStudy (aum):

We are trying to solve for k. 2 = k * (-1/2) To solve for k, we need to isolate k. On the right hand side, there is a 2 in the denominator. If we multiply both sides by 2, we can get rid of the 2 in the denominator: 2 * 2 = k * (-1/2) * 2 4 = k * (-1) 4 = -k multiply both sides by -1: -4 = k or k = -4 So y = -4x

OpenStudy (aum):

@megan_tanuis The problem states: "Find each constant variation. Then find the value of y when x= -0.3" So we have to find k in each case.

OpenStudy (anonymous):

But its telling you to find y. if you set it up like \[\frac{ x }{ y }\times \frac{ x }{ y }\] then cross multiply, you can get the answer that way too

OpenStudy (anonymous):

Unless I'm taking this the wrong way and you arent supposed to use the first set of numbers given with the others in 1-5

OpenStudy (anonymous):

Then yes, you're right if thats the case.

OpenStudy (aum):

Each problem has its own constant of variation. They want the "constant of variation" for each problem. Then they want the y value when x = -0.3 for each problem.

OpenStudy (anonymous):

@aum okay so for #2 \[y=\frac{ 2 }{ 3 }, x=0.3\] y=kx \[\frac{ 2 }{ 3}=k \times 0.3\] multiply both sides by 3? and i'll get \[3\times \frac{ 2 }{ 3 }=k \times0.3\times3\] \[9=k\times0.9\]?

OpenStudy (aum):

3 x 2/3 = 2 (not 9)

OpenStudy (anonymous):

oh okay.. so i'll end up having.. \[2=k\times0.9\] and multiply each side by 2?

OpenStudy (aum):

We want to isolate k. So we got to get rid of the 0.9. Divide both sides by 0.9 2/0.9 = k Normally it is not a good practice to have both decimal and fraction together. So get rid of the decimal by multiplying top and bottom by 10: k = 2 / 0.9 = 2*10 / (0.9 * 10) = 20/9

OpenStudy (aum):

So for problem #2, the constant of variation is 20/9 y = kx y = 20/9 * x Find y when x = -0.3 y = 20/9 * (-0.3) = -6/9 = -2/3

OpenStudy (anonymous):

so \[y=-\frac{ 2 }{ 3 }\] ?

OpenStudy (aum):

I just noticed you made a mistake in copying the problem about 6 replies before this. Problem #2 is: y= 2/3 when x= 0.2 (you had it as 0.3) y = kx 2/3 = k * 0.2 multiply by 3: 2 = 0.6 * k divide by 0.6 2/0.6 = k multiply top and bottom of fraction by 10 to get rid of the decimal: 20/6 = k simplify fraction: 10/3 = k or k = 10/3 is the constant of variation. "find the value of y when x= -0.3" y = kx y = 10/3 * x x = -0.3 y = 10/3 * (-0.3) = 10 * (-0.1) = -1

OpenStudy (aum):

3. y=7 when x= 2 y = kx 7 = k * 2 divide by 2: 7/2 = k or k = 3.5 is the constant of variation y = 3.5x "find the value of y when x= -0.3" y = 3.5 * (-0.3) = -1.05 You can finish the fourth problem the same way.